A spacecraft antenna is at an angle of 10◦


relative to the axis of the spacecraft. If the

spacecraft moves away from the earth at a speed of 0.70c, what is the angle of the antenna

as seen from the earth?


1
Expert's answer
2022-04-26T14:56:08-0400

"\theta'=\arctan{\frac{\tan\theta}{\sqrt{1-\frac{v^2}{c^2}}}}=14°."